# Perception-aware Sampling for Scatterplot Visualizations

### v1

@@ Abstract.
Visualizing data is often a crucial first step in data analytics workflows, but growing data sizes pose challenges due to computational and visual perception limitations. As a result, data analysts commonly down-sample their data and work with subsets. Deriving representative samples, however, remains a challenge. This paper focuses on scatterplots, a widely-used visualization type, and introduces a novel sampling objective—perception-awareness—aiming to improve sample efficacy by targeting humans’ perception of a visualization.
We make the following contributions: (1) We propose perception-augmented databases and design PAwS: a novel perception-aware sampling method for scatterplots that leverages saliency maps—a computer vision tool for predicting areas of attention focus in visualizations—and models perception-awareness via saliency, density, and coverage objectives. (2) We design ApproPAwS: a fast, perception-aware method for approximate visualizations, which exploits the fact that small visual perturbations are often imperceptible to humans. (3) We introduce the concept of perceptual similarity as a metric for sample quality, and present a novel method that compares saliency maps to measure it. (4) Our extensive experimental evaluation shows that our methods consistently outperform prior art in producing samples with high perceptual similarity, while ApproPAwS achieves up to 100× speed-ups with minimal loss in visual fidelity. Our user study shows that PAwS is often preferred by humans, validating our quantitative findings.
@@ 1. Introduction
Figure 1. The uniform random sample (b) of the original dataset (a) obscures the presence of the two smaller clusters. A sample of the same size, drawn with our perception-aware sampling method (c), better preserves key features of the original dataset, including trends, clusters, and outlier spread.
@@
Contributions and outline. To the best of our knowledge, our work is the first to explicitly model perception as an objective in sampling for scatterplot visualizations. We organize our contributions as follows.
@@
2. Perception-Aware Sampling: Overview and Background
@@ 2.1. Modeling perception through saliency
Figure 4. The DVS model (Matzen et al., 2017; Shin et al., 2022) derives a saliency map (b) of a scatterplot visualization (a). High saliency identifies areas that are likely to draw human focus. Human attention typically does not focus on the interior of dense areas, so the center of the map has low saliency. This heatmap can be projected on the data to assign saliency weights (c). Perception weights (d) augment saliency with density to indicate high perceptual significance for areas that are highly salient or highly dense.
Accounting for high-density areas. Saliency models are designed to represent human attention, which is an intuitive proxy for perception, with a one caveat: human focus is drawn on the perimeter of highly-dense elements, ignoring their interior. As a result, the saliency of a solid object is hardly indistinguishable from that of a hollow object (Figure 3); yet, most would agree that these shapes are perceived differently. Thus, we augment saliency models with density information, implementing the following intuition: an area of high saliency or high density should have a high perceptual value.
To estimate density, we employ non-parametric kernel density estimators (Scott, 2015; Silverman, 2018; Parzen, 1962). More formally, given a $d$-dimensional dataset $\mathcal{D}$ consisting of $n$ points and a density estimation model $\mathcal{F}$, the estimator $\mathcal{F}(\mathcal{D}):\mathbb{R}^{n\times d}\rightarrow\mathbb{R}^{n}$ computes a density estimate for each point in $\mathcal{D}$. Figure 4d shows the heatmap of perception weights computed over the data of Figure 4a; these combine the saliency weights (Figure 4c) with density information. We discuss the details of our approach in Section 3.
@@
Perception-aware sampling. We are now ready to provide a formal definition of the problem of perception-aware sampling:
@@ 2.2. Perception-aware sampling algorithm
Definition 2.1 seeks the sample of size $k$ that results in the most similar saliency map as that of the original data. However, since the saliency model and the SSIM measure (which is our perceptual similarity function) are blackbox components, the objective cannot be optimized directly. Iterative approaches may be possible, but the computational cost of the saliency model makes repeated calls to it impractical. Thus, we need to pursue heuristic methods for emulating the desired behavior.
The naive approach of using perception weights as sampling probabilities performs poorly. Intuitively, the reason is that the saliency of an area can be impacted by data in surrounding areas. Consider for example that outliers may draw human attention, but if the sample only contained outlier points, they would no longer be perceived as outliers and the attention patterns could change. Figure 5(b) shows a sample drawn from the data of Example 1.1, by using perception weights alone as sampling probabilities. We note that the result resembles uniform random sampling, and similarly fails to capture the small clusters. Based on this intuition, regardless of the perceptual model, perception weights cannot be used in a straightforward way to achieve perception-awareness. Our key insight is that, to avoid perceptual distortion in the sample, we need to prioritize areas of high perceptual significance, and provide good coverage of the overall data distribution. To achieve coverage, we turn to diverse data selection methods.
Max-Min diversification. Diversity and coverage are related objectives in data selection, and are well-studied by prior work (Drosou et al., 2017; Moumoulidou et al., 2021; Addanki et al., 2022; Borodin et al., 2017; Abbassi et al., 2013). The goal of data diversification is to select a sample such that the diversity across the selected points is maximized. Prior work has introduced a variety of diversification models; we focus on Max-Min, which is one of the most well-established and frequently-used models (Drosou et al., 2017).
Max-Min diversification, also known as farthest point sampling in the visualization and computer vision communities (Quadri et al., 2022; Eldar et al., 1997; Lin et al., 2022), selects points that are uniformly dispersed across the data space, and thus provides coverage and preserves the original data shape (Drosou and Pitoura, 2013; Wang et al., 2018; Moumoulidou et al., 2021). Similar to VAS, which also employs a coverage-based loss function, Max-Min sampling is susceptible to noise. However, Max-Min produces a better sample than VAS on the data of Example 1.1 (Figure 5(c)) and is also computationally more efficient. The objective of Max-Min is to maximize the diversity of a sample, defined as the minimum pairwise distance of the selected points.
@@
Figure 5. Perception weights for the dataset in Figure 1(a). (b) A weight-based sample, which selects points with probabilities proportional to their perception weights, obscures the outlier trend and the two smaller clusters. (c) Max-Min provides uniform data coverage but fails to reveal clusters and trends in the data that become apparent in the saliency map.
Perception-aware sampling. For Max-Min diversification, there exists a simple approximation algorithm, GMM, that relies on the farthest-first traversal heuristic (Ravi et al., 1994). GMM starts with a random point and keeps selecting the point whose minimum distance to previously selected points is maximized. We design a hybrid method based on the GMM algorithm that employs perception weights through a distance re-scaling mechanism. Our perception-aware sampling algorithm, PAwS, inherits the simplicity and intuition of Max-Min sampling, but re-scales the pairwise similarity scores of data points using the perception weights. This way, it achieves a good balance between coverage and prioritizing the selection of points in areas of high perceptual significance, organically adapting to different data settings. We describe PAwS in Section 3.
@@ 2.3. Approximate visualizations
In visual data exploration, analysts typically look for patterns, trends, and other insights, but they do not typically look at visualizations to read individual data points. In scatterplots in particular, humans are unlikely to be able to visually perceive small value perturbations in the data, which means that we can allow inaccuracies in the presented data, as long as the visualization preserves the overall trends (Healey and Enns, 2011). Our key insight is that we can generate an approximate visualization, that does not access or sample the real data at all. We achieve this through a novel perception-aware compression scheme, which works by fragmenting the visual canvas into rectangular regions with similar perception weights; considerations of area size and data density allow for adjusting the level of compression. A modified PAwS sampling selects boxes to sample from, and then draws points uniformly at random within a box.
Figure 6(c) demonstrates the approximate visualization generated by our ApproPAwS strategy. It displays the same number of points as our PAwS sample (Figure 6(b)), but these points were not drawn from the original data, but rather, they were “constructed” through our perception-aware compression scheme. Yet, a human would be extremely unlikely to recognize that this is not an actual sample from the original data (Figure 6(a)). Our ApproPAwS strategy closely emulates the behavior of PAwS, but because it works on the compressed representation and does not access the data directly, it can result in significant runtime gains in large datasets. We discuss the details of this method in Section 4.
@@ 2.4. Scope and practical considerations
Our goal in this paper is to establish that objectives targeting human perception (modeled via saliency) achieve more effective data samples compared to state-of-the-art methods. We provide proof-of-concept implementation and demonstrate the effectiveness of perception-aware sampling through extensive quantitative experiments and a user study (Section 6). It is not our goal to evaluate how well state-of-the-art saliency models capture visual attention. This has already been established in prior work (Bylinskii et al., 2016, 2018; Matzen et al., 2017; Shin et al., 2022; Itti and Koch, 2001), and we use these state-of-the-art tools as blackboxes.
@@ 3.1. Perception-augmented database
As we described in Section 2, our perception model is build over saliency maps. However, producing a suitable saliency map is more than simply invoking a saliency model. Human perception, and by extension saliency, is affected by the visual features of a visualization (e.g., point size, opacity, canvas aspect ratio) (Micallef et al., 2017; Kwon et al., 2017). In Figure 8(a), we render the same dataset with varying configurations for the point size (2–16) and opacity (0.1–1), with an aspect ratio of 1.25:1. Figure 8(b) shows the saliency map that corresponds to each rendering. We observe that for small point sizes and opacity, the visualization highlights areas of higher population density, and the corresponding saliency maps identify these areas as more salient. For large point sizes and opacity, more areas of the visualization appear dense, and the attention focus tends to shift towards the outskirts and outliers.
Selecting a good configuration for a scatterplot is non-trivial and it is an active problem in visualization research (Micallef et al., 2017; Quadri et al., 2022; Kwon et al., 2017; Quadri and Rosen, 2020). Our observations in Figure 8 further demonstrate that different configurations emphasize different aspects of the data. To better preserve these aspects, we follow a stimuli-driven approach to generate saliency maps across different configurations and then aggregate them into a single heatmap (Figure 8(c)). Different aggregation options are possible; in our implementation, we use the maximum saliency value of each pixel across all configurations, as we found it to be more robust and better at preserving different data trends.
@@
Figure 8. Visual features affect visualizations and saliency. (a) Data visualizations for different point sizes and opacity configurations. (b) Saliency maps produced for each data visualization. (c) Aggregate saliency over the maps in (b).
Deriving perception weights. The perception-augmented augmented database translates the aggregate saliency map into saliency weights. The saliency map is a 2D matrix assigning saliency values to pixels, which are projected onto data points in the visualization to derive saliency weights (Figure 4). The database stores aggregate saliency maps with a small storage overhead, and saliency weights are computed on the fly with negligible impact on runtime performance. [^2] Saliency weights could be stored in the perception-augmented database, but while storage overhead per point is constant, the total storage overhead would grow with the size of the data. To derive perception weights (Section 2), we augment saliency with density. We estimate density using non-parametric kernel density estimators (Scott, 2015; Silverman, 2018; Parzen, 1962), assigning higher density weights to points in highly dense areas. Then, for a data point $p$, with a saliency weight $q_{1}$ and a density weight $q_{2}$, the perception weight is equal to $w_{p}=\textup{max}(q_{1},\gamma\cdot q_{2})$, where $\gamma\in[0,1]$ balances the influence of density.
Density weights are stored as an additional data attribute, while perception weights are computed on the fly. Using the maximum operator ensures that points are prioritized while sampling if located in highly salient or highly dense areas. To adaptively set $\gamma$, we measure the variance of the density weights: when variance is low, density adds little information as all areas have similar density ($\gamma$ approaches zero); when variance is high, density plays a larger role. In our experiments, we map the density variance of a dataset into a sigmoid function to obtain a $\gamma\in[0,1]$, assigning $\gamma$ values closer to one for datasets with larger density variance. Figure 4 shows an example of deriving perception weights.
@@ 3.2. Perception-aware data selection
Our goal is to select a sample $\mathcal{D^{\prime}}\subseteq\mathcal{D}$ of size $k$ that maximizes the perceptual similarity of $\mathcal{D^{\prime}}$ to the original dataset $\mathcal{D}$, i.e., the similarity of the saliency maps of $\mathcal{D^{\prime}}$ and $\mathcal{D}$ should be maximized. Given the blackbox nature of saliency models, PAwS follows a heuristic approach that balances coverage objectives with prioritizing the selection of points with high perception weights.
PAwS adapts the greedy selection strategy in Max-Min sampling: It initially selects a random point, and proceeds to augment the sample $\mathcal{D^{\prime}}$ greedily, by adding the point $p^{*}$ with the maximum score $s_{p^{*}}=w_{p^{*}}\cdot\underset{x\in\mathcal{D^{\prime}}}{\text{min}}\ d(p^{*},x)$. This score is defined as the product of the distance of the point to the existing sample and its perception weight. Thus, the objective penalizes points that have low perception weights, as well as points that are too close to the existing sample. This strategy is analogous to data augmentation techniques in computer vision that balance saliency and coverage (Lee et al., 2022).
Running time. Algorithm 1 has $O(kn)$ running time, if the implementation maintains the pairwise distances of each point to the sample ($\min_{x\in\mathcal{D^{\prime}}}d(y,x)$ in line 5) by only computing distances to each newly added point. Similar implementations have been used in the literature (Wang et al., 2018; Kleindessner et al., 2019).
@@
$\mathcal{D}$: Dataset of $n$ point
@@ 4. Approximate Visualizations
Figure 9 demonstrates the high-level workflow of our approximate visualization module. The perception-augmented DB uses a quad-tree partitioning scheme to produce a compressed representation of the data based on the perception model. Intuitively, the compressed representation fragments the canvas area into bounding boxes of different sizes, so that the data that falls within each box is somewhat evenly distributed and has similar perception weights. Our adapted sampling algorithm, ApproPAwS (Algorithm 2), generates a set of representative points for each box, drawn uniformly at random within the box area. It then greedily selects a point at a time, based on a combination of perception and coverage objectives (like PAwS), replacing the point with another random one in the same box. ApproPAwS achieves runtime gains simply on the premise of working over a small dataset of randomly-drawn points, so these gains are more pronounced the larger the original data is.
Figure 9. The perception-augmented DB can store a compressed representation of the data using a quad-tree partitioning structure. Each cell conforms to uniformity requirements for the data distribution and perception weights, which control the compression level. ApproPAwS (Algorithm 2) can then generate a sample directly from the perception-aware compression, resulting in an approximate visualization.

### v3

@@ Abstract.
Visualizing data is often central in data analytics workflows, but growing data sizes pose challenges due to computational and visual perception limitations. As a result, data analysts commonly down-sample their data and work with subsets. Deriving representative samples, however, remains a challenge. This paper focuses on scatterplots, a widely-used visualization type, and introduces a novel sampling objective—perception-awareness—aiming to improve sample efficacy by targeting how humans perceive a visualization.
We make the following contributions: (1) We propose perception-augmented databases and design PAwS: a novel perception-aware sampling method for scatterplots that leverages saliency maps—a computer vision tool for predicting areas of attention focus in visualizations—and models perception-awareness via saliency, density, and coverage objectives. (2) We design ApproPAwS: a fast, perception-aware method for approximate visualizations, which exploits the fact that small visual perturbations are often imperceptible to humans. (3) We introduce the concept of perceptual similarity as a metric for sample quality, and present a novel method that compares saliency maps to measure it. (4) Our extensive experimental evaluation shows that our methods consistently outperform prior art in producing samples with high perceptual similarity, while ApproPAwS achieves up to 100x speed-ups with minimal loss in visual fidelity. Our user study shows that PAwS is often preferred by humans, validating our quantitative findings.
@@ 1. Introduction
Figure 1. A random sample (b) of the original dataset (a) obscures the two smaller clusters. A sample of the same size, drawn with our perception-aware sampling method (c), better preserves key features of the original dataset, including trends, clusters, and outlier spread.
@@
Contributions and outline. To the best of our knowledge, our work is the first to explicitly model perception as an objective in sampling for scatterplot visualizations. In this paper:
@@
2. Overview and Background
@@ 2.1. Modeling perception through saliency
Figure 4. The DVS model (Matzen et al., 2017; Shin et al., 2022) derives a saliency map (b) of a scatterplot visualization (a). Human attention typically does not focus on the interior of dense areas, so the center of the map has low saliency. This heatmap can be projected on the data to assign saliency weights (c). Perception weights (d) augment saliency with density to indicate high perceptual significance for areas that are highly salient or highly dense.
Accounting for high-density areas. Saliency models are designed to represent human attention, which is an intuitive proxy for perception, with one caveat: human focus is drawn on the perimeter of highly-dense elements, ignoring their interior. As a result, the saliency of a solid object is hardly indistinguishable from that of a hollow object (Figure 3); yet, most would agree that these shapes are perceived differently. Thus, we augment saliency models with density information, implementing the following intuition: an area of high saliency or high density should have a high perceptual value. Figure 4d shows the heatmap of perception weights computed over the data of Figure 4a; these combine the saliency weights (Figure 4c) with density information. We discuss the details in Section 3.
@@
We are now ready to formalize our problem definition.
@@ 2.2. Perception-aware sampling algorithm
Definition 2.1 seeks the sample of size $k$ that results in the most similar saliency map as that of the original data. However, since the saliency model and the SSIM measure (which is our perceptual similarity function) are blackbox components, the objective cannot be optimized directly. Iterative approaches may be possible, but the computational cost of the saliency model makes repeated calls to it impractical. Thus, we explore heuristic methods.
The naive approach of using perception weights as sampling probabilities performs poorly (Figure 5(b)). Intuitively, the reason is that the saliency of an area can be impacted by data in surrounding areas. For example, outliers may draw human attention, but if the sample only contained outlier points, they would no longer be perceived as outliers and the attention patterns could change. Based on this intuition, perception weights should be used in unison with coverage objectives to reduce perceptual distortion in the sample.
Max-Min diversification. Diversity and coverage are related objectives in data selection, and are well-studied by prior work (Drosou et al., 2017; Moumoulidou et al., 2021; Addanki et al., 2022; Borodin et al., 2017; Abbassi et al., 2013). Prior work has introduced a variety of diversification models; we focus on Max-Min, which is one of the most well-established and frequently-used models (Drosou et al., 2017).
Max-Min diversification, also known as farthest point sampling in the visualization and computer vision communities (Quadri et al., 2022; Eldar et al., 1997; Lin et al., 2022), selects points that are uniformly dispersed across the data space, and thus provides coverage and preserves the original data shape (Drosou and Pitoura, 2013; Wang et al., 2018; Moumoulidou et al., 2021). The objective of Max-Min is to maximize the diversity of a sample, defined as the minimum pairwise distance of sample points. As with other coverage-based objectives, Max-Min is susceptible to noise (Figure 5(c)), but it is computationally more efficient than VAS.
@@
Figure 5. Perception weights for the dataset in Figure 1(a). (b) A weight-based sample, which selects points with probabilities proportional to their perception weights, obscures the outlier trend and the two smaller clusters. (c) Max-Min provides coverage but fails to reveal clusters and trends in the data that are apparent in the saliency map.
@@
Perception-awareness and coverage. GMM, a simple approximation algorithm for Max-Min diversification, relies on the farthest-first traversal heuristic (Ravi et al., 1994). GMM starts with a random point and keeps selecting the point whose minimum distance to previously selected points is maximized. We design a hybrid method based on the GMM algorithm that employs perception weights through a distance re-scaling mechanism. Our perception-aware sampling algorithm, PAwS, inherits the simplicity and intuition of Max-Min sampling, but re-scales the pairwise similarity scores of data points using the perception weights, thus balancing coverage and prioritizing points with high perception value. We describe PAwS in Section 3.
@@ 2.3. Approximate visualizations
In visual data exploration, analysts typically look for patterns, trends, and other insights, but they do not typically look at visualizations to read individual data points. In scatterplots in particular, humans are unlikely to visually perceive small value perturbations in the data, which means that we can allow inaccuracies in the presented data, as long as the visualization preserves the overall trends (Healey and Enns, 2011). Our key insight is that we can generate an approximate visualization, that does not access or sample the real data at all. We achieve this through a novel perception-aware compression scheme, which works by fragmenting the visual canvas into rectangular regions with similar perception weights; considerations of area size and data density allow for adjusting the level of compression.
Figure 6(c) demonstrates the approximate visualization generated by our ApproPAwS strategy. It displays the same number of points as our PAwS sample (Figure 6(b)), but these points were not drawn from the original data; rather, ApproPAwS modies PAwS to select boxes from the compressed representation to sample from, and then draws points uniformly at random within a box. Yet, a human would be extremely unlikely to recognize that this is not an actual sample from the original data (Figure 6(a)). Our ApproPAwS strategy closely emulates the behavior of PAwS, but because it works on the compressed representation and does not access the data directly, it can result in significant runtime gains in large datasets. We discuss the details of this method in Section 4.
@@ 2.4. Scope and practical considerations
Our goal in this paper is to establish that objectives targeting human perception (modeled via saliency) achieve more effective data samples compared to state-of-the-art sampling methods. We provide proof-of-concept implementation and demonstrate the effectiveness of perception-aware sampling through extensive quantitative experiments and a user study (Section 6). It is not our goal to evaluate how well state-of-the-art saliency models capture visual attention. This has already been established in prior work (Bylinskii et al., 2016, 2018; Matzen et al., 2017; Shin et al., 2022; Itti and Koch, 2001), and we use these state-of-the-art tools as blackboxes.
@@ 3.1. Perception-augmented database
Saliency aggregation. Different configurations of visual features (e.g., point size, opacity, canvas aspect ratio) can affect the appearance of a visualization and thus its saliency (Micallef et al., 2017; Kwon et al., 2017). Selecting a good configuration for a scatterplot is non-trivial and it is an active problem in visualization research (Micallef et al., 2017; Quadri et al., 2022; Kwon et al., 2017; Quadri and Rosen, 2020). In Figure 8(a), we render the same dataset with varying configurations for the point size (2–16) and opacity (0.1–1), with an aspect ratio of 1.25:1. Figure 8(b) shows the saliency map that corresponds to each rendering. We observe that for small point sizes and opacity, the visualization highlights areas of higher population density, and the corresponding saliency maps identify these areas as more salient. For large point sizes and opacity, more areas of the visualization appear dense, and the attention focus tends to shift towards the outskirts and outliers. Focusing on a particular configuration and saliency would therefore potentially miss important perspectives in the data that become salient under different configurations. For a more robust perception model, we aggregate saliency information across a broad range of configurations (Figure 8(c)). In our implementation, we use the maximum saliency value of each pixel across all configurations, but other aggregation options are possible.
@@
Figure 8. The same dataset will look different with different point size and opacity configurations (a), which will translate to different saliency maps (b). Our model aggregates the saliency across varied stimuli to capture all these perspectives (c).
Deriving perception weights. The perception-augmented database stores two pieces of information as meta data: the aggregate saliency map (a 2D matrix assigning saliency values to pixels) for the target visualization, and a density score $q_{d}$ for each data point. The system derives a per-tuple saliency score $q_{s}$ on the fly (with negligible overhead), by assigning to each datapoint the corresponding pixel value. The density score is precomputed using non-parametric kernel density estimators (Scott, 2015; Silverman, 2018; Parzen, 1962), assigning higher density weights to points in highly dense areas. The perception-augmented DB then derives a perception weight for each point as $w_{p}=\max(q_{s},\gamma\cdot q_{d})$, where $\gamma\in[0,1]$ controls the influence of density. The system adaptively sets $\gamma$ based on the variance of density weights: when variance is low, density adds little information as all areas have similar density ($\gamma$ approaches zero); when variance is high, density plays a larger role. In our experiments, we map the density variance of a dataset into a sigmoid function, assigning $\gamma$ values closer to one for datasets with larger density variance.
@@ 3.2. Perception-aware data selection
Our goal is to select a sample $\mathcal{D^{\prime}}\subseteq\mathcal{D}$ of size $k$ that maximizes the perceptual similarity of $\mathcal{D^{\prime}}$ to the original dataset $\mathcal{D}$, i.e., the similarity of the saliency maps of $\mathcal{D^{\prime}}$ and $\mathcal{D}$ should be maximized. We use coverage and perception scores as proxies towards this objective. PAwS follows a greedy strategy, adapted from Max-Min sampling: It initially selects a random point, and proceeds to augment the sample $\mathcal{D^{\prime}}$ greedily, by adding the point $p^{*}$ with the largest multiplicative score of its perception weight ($w_{p^{*}}$) and distance from the current sample ($\min_{x\in\mathcal{D^{\prime}}}d(p^{*},x)$). Thus, the objective penalizes points that have low perception weights, or points that are too close to the existing sample. This strategy is analogous to data augmentation techniques in computer vision that balance saliency and coverage (Lee et al., 2022).
Running time. To maintain the pairwise distances of each point to the sample (line 5), we only need to recompute distances to each newly-added point, resulting in $O(kn)$ running time. Similar implementations have been used in the literature (Wang et al., 2018; Kleindessner et al., 2019).
@@
$\mathcal{D}$: Dataset of $n$ points
@@ 4. Approximate Visualizations
Figure 9 demonstrates the high-level workflow of our approximate visualization module. The perception-augmented DB uses a quad-tree partitioning scheme to produce a compressed representation of the data based on the perception model. Roughly, the compressed representation fragments the canvas area into bounding boxes of different sizes, so that the data that falls within each box is somewhat evenly distributed and has similar perception weights. Our adapted sampling algorithm, ApproPAwS (Algorithm 2), generates a set of representative points for each box, drawn uniformly at random within the box area. It then greedily selects a point at a time, based on a combination of perception and coverage objectives (like PAwS), replacing the point with another random one in the same box. ApproPAwS achieves runtime gains simply on the premise of working over a small dataset of randomly-drawn points, so these gains are more pronounced the larger the original data is.
Figure 9. The perception-augmented DB can store a compressed representation of the data using a quad-tree partitioning structure. Each cell conforms to uniformity requirements for the data distribution and perception weights, which control the compression level. ApproPAwS (Algorithm 2) can derive a sample directly from the compressed representation, resulting in an approximate visualization.